. Suppose that for a particular economy and period, investment was equal to 200, government expenditure was equal to 100, net taxes were fixed at 150, and consumption (C) was given by the consumption function C = 20 + 0.6YD where YD is disposable income and Y is GDP. a. What is the level of equilibrium income (Y) ? b. What is the value of the government expenditure multiplier (Y>G) ? Of the tax multiplier (Y>T)? c. Suppose the investment declined by 100 units to a level of 100. What will be the new level of equilibrium income?
1
Expert's answer
2020-07-17T12:41:59-0400

"a) Y = C + I + G + (X \u2013 M)= 20 + 0.6(Y-150)+200+100\n = 20+0.6Y-90+200+100" "Y=\frac{230}{0.4}=575"



b) Government expenditure multiplier="\frac{575}{100}=5.75"

Tax multiplier= "\frac{575}{150}=3.833"



c) "Y=\frac{130}{0.4}=325"